What is the tripal point of water on a Fahrenheit scale? What is the absolute zero on this
scale 

Grade:11

1 Answers

Arun
5 years ago
The triple point of pure water is at 0.01°C (273.16K, 32.01°F).
 
absolute zero = – 273oc
 
hence f – 32/9 = c/5
 
f- 32 = -9* 54.6 = – 491.4
 
f = – 459.4 farenhite